# The interplay of the parametric instability and MRI in oscillatory shear flows

### v1

@@ 4.3 Equilibrium sloshing
Figure 7: The sloshing equilibrium established for our $\psi_{\rm f}=0.3$, hydrodynamic run. Panel (a): meridional cross-section of $\{u_{x}\}_{y}$ at $t=60T$. Panel (b): $\{u_{x}\}_{xy}$ profile as a function of $z$ for different oscillation phases over the orbital period $4140<t/T<$ 41. Panel (c): likewise for $\{u_{y}-u_{\rm K}\}_{xy}$. Panel (d): the best fit sinusoidal profile for the radial sloshing $A(t)$ (black line) and azimuthal sloshing $B(t)$ (red line). Panel (e): polar plot of the Fourier amplitudes and phases of the extracted $A(t)$ (black) and $B(t)$ (red) oscillations, corresponding to the dynamical frequency.
@@
After $t/T\sim 10$ the simulation broadly equilibrates into a regular sloshing back and forth, as indicated by the leveling out of the epicylic energy in panel (a) of Fig. 5. We extend this run up until $t/T=60$ to ensure we have attained a quasi-steady state. The structure of this flow is visualized in Fig. 7. In panel (a) we plot the density-weighted, y-averaged radial velocity $\{u_{x}\}_{y}$ at the final snapshot $t=60T$, across the meridional $x$-$z$ plane. Here we clearly see the established sloshing flow, which is approximately odd about the midplane. However, we also recognize that this flow is distinct from the smooth, laminar picture presented in section 2. Indeed, the nonlinear manifestation of the PI distorts the interface between the sloshing layers as the strong vertical flows, discussed in section 4.2, lead to a corrugation pattern. The magnitude of the sloshing flow is also notably reduced compared to the supersonic velocities attained during the resonant evolution, before the PI kicks in. In fact, we can obtain more quantitative information by performing a horizontal, density-weighted average of the non-Keplerian velocity field. We plot the $z$-profiles of $\{u_{x}\}_{xy}$ and $\{u_{y}-u_{K}\}_{xy}$ in panels (b) and (c) respectively. The colored lines show the sloshing profiles at different times during the orbital forcing phase (here plotted for the orbit $4140<t/T<$ 41). These nicely exemplify the sinusoidal and oscillatory behavior of the averaged sloshing motions – in keeping with a simplified, viscous, laminar model. For each snapshot, we use a least-squares procedure to fit the profiles with a sinusoidal function $\propto\sin(k_{\rm b}z)$ in order to extract the sloshing amplitudes $A(t)$ and $B(t)$. These are plotted as a function of time in panel (d) where the radial sloshing $A(t)$ and azimuthal sloshing $B(t)$ are shown as the black and red lines respectively. These exhibit steady amplitudes and maintain a coherent, anti-phased relationship. We perform a Fourier analysis on these time series and find the dominant frequency matches the orbital frequency, as expected. The Fourier component at this frequency measures the signal amplitude and phase as shown in polar plot in panel (e). Writing $A(t)=\mathrm{Re}\{|A|\exp[i(\arg(A)+\Omega_{0}t)]\}$ and likewise for $B(t)$, we find $|A|=0.185$ and $\mathrm{arg}(A)=-0.76$, whilst $|B|=0.054$ and $\mathrm{arg}(B)=2.49$ – yielding an amplitude ratio of $3.43$ and a phase difference of $1.03\pi$.
@@
We can test how well this behavior is captured by the laminar, viscous model presented in section 2. Specifically, this model predicts an equilibrium described by equations (15)-(16) which encapsulate the amplitude and phase relationship of the $x$ and $y$ sloshing motions. Since we have two free parameters, $\nu_{xz}$ and $\nu_{yz}$, we will try to optimize these by minimizing the cost function,
@@
The results are shown in Fig. 8 wherein the contour plot visualizes level sets of the cost function. This is minimized at the location of the red star for which $(\nu_{xz},\nu_{yz})=(0.0256,0.0354)$. This yields $|A|_{\rm vis}=0.185$ and $|B|_{\rm vis}=0.054$ in accordance with the amplitudes extracted from the simulation. However, the resulting model phase difference $\textrm{arg}(B_{\rm vis})-\textrm{arg}(A_{\rm vis})=0.802\pi$ is clearly different. This disparity is indicative of the shortcomings of a viscous, shearing prescription in capturing the full details of a turbulent, stress relationship. Even more simply, if we restrict our attention to isotropic viscosities along the white dashed line, we find that the cost function is minimized for $\nu_{xy}=\nu_{yz}=0.0295$ at the position of the blue star. [^4] Notice that the value of $\nu$ extracted here is too large to apply the simpler formula given by equation (17). This simplified isotropic model then predicts $|A|_{\rm vis}=0.171$ and $|B|_{\rm vis}=0.056$, with $\textrm{arg}(B_{\rm vis})-\textrm{arg}(A_{\rm vis})=0.774\pi$, which is a marginally worse fit to the simulation results, compared with the anisotropic model considered above. This is unsurprising since the isotropic assumption reduces the available degrees of freedom for fitting the data. Moreover, as we will see in the following section, these purely viscous prescriptions will be supplanted by more flexible stress-strain models, which perform even better at fitting the data.
@@ 4.4 Reynolds Stresses
akin to equations (8)-(9) of Torkelsson et al. (2000) but with the additional forcing term. Here we work with the velocity components $\mathbf{v}=\mathbf{u}-\mathbf{u}_{\rm K}$, relative to the underlying Keplerian flow. The horizontal averaging extracts the coherent sloshing motion which we will denote $\mathbf{v}_{\rm w}\equiv\langle\rho\mathbf{v}\rangle_{xy}/\bar{\rho}=\{\mathbf{v}\}_{xy}$, where the density average $\bar{\rho}=\langle\rho\rangle_{xy}$. Note that we can identify the in-plane, averaged components with the sloshing amplitudes $A$ and $B$ as defined in equations (6)-(7). Therefore, we decompose the non-Keplerian velocity into a sloshing part and a turbulent residual $\mathbf{u}^{\prime}$ such that $\mathbf{v}=\mathbf{v}_{\rm w}+\mathbf{u}^{\prime}$, where $\{\mathbf{u}^{\prime}\}_{xy}=0$. Propagating this splitting through equations (25)-(26) we find
@@
Figure 9: The rate-of-strain and stress relationship for our fiducial hydro run. The upper panels show the phase-folded, vertical profiles for: (a) $v_{x,{\rm w}}$, (b) $v_{y,{\rm w}}$, (c) $\mathcal{R}_{xz}$ and (d) $\mathcal{R}_{yz}$. In panel (e) we extract the best-fitting, sloshing amplitudes $A(t)$ (black) and $B(t)$ (red) across the orbital phase. In panel (f) we similarly extract the best-fitting stress amplitudes, $R_{xz}(t)$ (black-dashed) and $R_{yz}(t)$ (red-dashed).
[…]
@@
This model suggests that the Reynolds stresses should be $\pi$ out of phase with the sloshing flows. In fact, we see that $\arg(R_{xz})-\arg(A)=1.03\pi$ and $\arg(R_{yz})-\arg(B)=-1.19\pi$. Although these are relatively consistent with the viscous picture, there are clear discrepancies. This reemphasizes the results of section 4.3, wherein the best viscous fit for the equilibrium sloshing could readily capture the sloshing amplitudes, but not their phase relationship. Therefore, we are encouraged to relax the purely viscous prescription and instead appeal to a more flexible closure model for the turbulence. Indeed, by assuming that $\nu$ is real we are implicitly stating that the stress responds instantaneously to the rate-of-strain. More realistically, the turbulent transport depends on the history of the shearing motion – in other words there is some relaxation or delay timescale between the shear and the stress response. Such time delays are also reminiscent of local simulations of the MRI, wherein small lags are found between energy injection rates, via the boundary stresses, and the thermalization rate, as channel flows take some finite time to subsequently decay via parasitic instabilities (see Sano & Inutsuka, 2001; Gardiner & Stone, 2005). This suggests the use of a viscoelastic model for which the stress and strain rate can have an arbitrary phase relationship (e.g. Ogilvie & Proctor, 2003; Ogilvie & Lesur, 2012). We therefore promote $\nu$ to being a complex quantity. Solving equation (30) for the viscoelastic components gives $\nu_{xz}=0.015+0.001i=0.015e^{0.09i}$ and $\nu_{yz}=0.043-0.030i=0.053e^{-0.60i}$. With this complex viscosity in hand, we can also generalize the predictions of the steady state, laminar sloshing given by equations (15)-(16). Inserting our new results for $\nu_{xz}$ and $\nu_{yz}$ we find that $|A|_{\rm vis}=0.186$ and $|B|_{\rm vis}=0.054$ with a phase difference of $\arg(B_{\rm vis})-\arg(A_{\rm vis})=1.03\pi$, almost in perfect agreement with the values extracted from the simulation. This should be expected given the added flexibility of the viscoelastic model, where each coefficient $\nu_{iz}$ provides two degrees of freedom, which can then fit two amplitudes and two phases.
@@ 5 MHD results
Having investigated the steady sloshing response for the hydrodynamic case we now wish to extend our setup to include magnetic fields. This will facilitate a complicated interplay between the MRI and PI which jointly determines the final sloshing state equilibrium. In section 5.1 we will characterize the base MRI state before the addition of sloshing flows. Then in section 5.2 we will examine how the growth of the PI proceeds in the presence of MRI turbulence. Next, in section 5.3 we will investigate the resulting equilibrium sloshing state and quantify the vertical magnetic and hydrodynamic stresses at play. Finally, in section 5.4 we will interrogate the dependence of our results on the forcing amplitude, comparing both the hydro and magnetized regimes.
@@ 5.1 Characterizing the base MRI
[…]
Figure 10: Characterization of our zero-net-flux, base MRI state. Panel (a): plots $\log_{10}(\beta_{\rm m})$ through the meridional slice $y=0$, showing field amplification and fine turbulent structure. Panel (b): growth and saturation of the horizontal Maxwell stresses $\alpha_{\rm M}$ (red), Reynolds stresses $\alpha_{\rm R}$ (blue) and net stress (black), all normalized by the box-averaged pressure. Panel (c): the density-weighted PDF ($f_{\rm PDF}$) for the velocity fluctuations $u_{x}^{\prime}$ (black), $u_{y}^{\prime}$ (red) and $u_{z}^{\prime}$ (blue), averaged over simulation snapshots between $10040<t/T<$ 100 in intervals of 10.
Representative summary diagnostics, characterizing the base MRI state, are shown in Fig. 10. In panel (a) we plot $\log_{10}(\beta_{\rm m})$ through the $y=0$ slice at $t=100T$, once the MRI has saturated. Here, we clearly observe the presence of turbulent structure over a range of scales. Furthermore, amplification of the magnetic field leads to eddies in which $\beta_{\rm m}$ is of order unity. The strength of the MRI is typically quantified in terms of the non-dimensionalized, horizontal accretion stresses it induces. These consist of the hydrodynamic Reynolds stresses
@@
Here we have normalized the stresses against the box-averaged pressure $\langle\rho c^{2}\rangle=\rho_{0}c^{2}=p_{0}=1$ and also the shear coefficient $q$, so that $\alpha$ can be identified with the Shakura-Sunyaev viscosity parameter acting on the horizontal shear flow. These average stresses are plotted as a function of time in panel (b) where $\alpha_{\rm R}$ (red line) and $\alpha_{\rm M}$ (blue line) are combined into the net stress (black line). Here we see a steep initial rise phase as the MRI grows over tens of dynamical orbits. As the instability saturates, the stresses turn over before leveling off beyond $t=40T$. The stresses are averaged over the gray shaded interval between $10040<t/T<$ 100, yielding typical values for $\alpha_{\rm M}=0.0069$ and $\alpha_{\rm R}=0.0014$. As per usual, the Maxwell stresses dominate over the hydrodynamic effects. In the final panel (c), we plot the density-weighted, normalized histogram ($f_{\rm PDF}$) for the velocity fluctuations, averaged over snapshots at $t/T\in[40,100]$ in intervals of 10. Here we see that the horizontal fluctuations tend to have a wider variance, with characteristic standard deviations $\sigma_{x}=0.085$ and $\sigma_{y}=0.079$, versus $\sigma_{z}=0.056$ for the vertical velocity. For a strong sloshing signal to visibly emerge in the forced simulations, it should manifest on velocities larger than these. However, even small sloshing amplitudes, disguised by the turbulence, can be extracted in an average sense.

### v2

@@ 4.3 Equilibrium Sloshing
Figure 7: Sloshing equilibrium established for our $\psi_{\rm f}=0.3$, hydrodynamic run. Panel (a): meridional cross section of $\{u_{x}\}_{y}$ at $t=60T$. Panel (b): $\{u_{x}\}_{xy}$ profile as a function of $z$ for different oscillation phases over the orbital period $40<t/T<41$. Panel (c): likewise for $\{u_{y}-u_{\rm K}\}_{xy}$. Panel (d): best-fit sinusoidal profile for the radial sloshing $A(t)$ (black line) and azimuthal sloshing $B(t)$ (red line). Panel (e): polar plot of the Fourier amplitudes and phases of the extracted $A(t)$ (black) and $B(t)$ (red) oscillations, corresponding to the dynamical frequency.
@@
After $t/T\sim 10$, the simulation broadly equilibrates into a regular sloshing back and forth, as indicated by the leveling off of the epicylic energy in panel (a) of Fig. 5. We extend this run until $t/T=60$ to ensure we have attained a quasi-steady state. The structure of this flow is visualized in Fig. 7. In panel (a), we plot the density-weighted, $y$-averaged radial velocity $\{u_{x}\}_{y}$ at the final snapshot $t=60T$, across the meridional xxzz plane. Here, we clearly see the established sloshing flow, which is approximately odd about the midplane. However, we also recognize that this flow is distinct from the smooth, laminar picture presented in Section 2. Indeed, the nonlinear manifestation of the PI distorts the interface between the sloshing layers as the strong vertical flows, discussed in Section 4.2, lead to a corrugation pattern. The magnitude of the sloshing flow is also notably reduced compared to the supersonic velocities attained during the resonant evolution, before the PI kicks in. In fact, we can obtain more quantitative information by performing a horizontal, density-weighted average of the non-Keplerian velocity field. We plot the $z$-profiles of $\{u_{x}\}_{xy}$ and $\{u_{y}-u_{K}\}_{xy}$ in panels (b) and (c), respectively. The colored lines show the sloshing profiles at different times during the orbital forcing phase (here plotted for the orbit $40<t/T<41$). These nicely exemplify the sinusoidal and oscillatory behavior of the averaged sloshing motions – in keeping with a simplified, viscous, laminar model. For each snapshot, we use a least-squares procedure to fit the profiles with a sinusoidal function $\propto\sin(k_{\rm b}z)$ in order to extract the sloshing amplitudes $A(t)$ and $B(t)$. These are plotted as a function of time in panel (d), where the radial sloshing $A(t)$ and azimuthal sloshing $B(t)$ are shown as the black and red lines, respectively. These exhibit steady amplitudes and maintain a coherent, antiphased relationship. We perform a Fourier analysis on these time series and find the dominant frequency matches the orbital frequency, as expected. The Fourier component at this frequency measures the signal amplitude and phase as shown in the polar plot in panel (e). Writing $A(t)=\mathrm{Re}\{|A|\exp[i(\arg(A)+\Omega_{0}t)]\}$ and likewise for $B(t)$, we find $|A|=0.185$ and $\mathrm{arg}(A)=-0.76$, whilst $|B|=0.054$ and $\mathrm{arg}(B)=2.49$ – yielding an amplitude ratio of $3.43$ and a phase difference of $1.03\pi$.
@@
We can test how well this behavior is captured by the laminar, viscous model presented in Section 2. In particular, this model predicts an equilibrium described by equations (15)-(16), which encapsulate the amplitude and phase relationship of the $x$ and $y$ sloshing motions. Since we have two free parameters, $\nu_{xz}$ and $\nu_{yz}$, we will try to optimize these by minimizing the cost function
@@
The results are shown in Fig. 8, wherein the contour plot visualizes level sets of the cost function. This is minimized at the location of the red star for which $(\nu_{xz},\nu_{yz})=(0.0256,0.0354)$. This yields $|A|_{\rm vis}=0.185$ and $|B|_{\rm vis}=0.054$ in accordance with the amplitudes extracted from the simulation. However, the resulting model phase difference $\textrm{arg}(B_{\rm vis})-\textrm{arg}(A_{\rm vis})=0.802\pi$ is clearly different. This disparity is indicative of the shortcomings of a viscous, shearing prescription in capturing the full details of a turbulent, stress relationship. Even more simply, if we restrict our attention to isotropic viscosities along the white dashed line, we find that the cost function is minimized for $\nu_{xy}=\nu_{yz}=0.0295$ at the position of the blue star. [^4] Notice that the value of $\nu$ extracted here is too large to apply the simpler formula given by equation (17). This simplified isotropic model then predicts $|A|_{\rm vis}=0.171$ and $|B|_{\rm vis}=0.056$, with $\textrm{arg}(B_{\rm vis})-\textrm{arg}(A_{\rm vis})=0.774\pi$, which is a marginally worse fit to the simulation results, compared with the anisotropic model considered above. This is unsurprising since the isotropic assumption reduces the available degrees of freedom for fitting the data. Moreover, as we will see in the following section, these purely viscous prescriptions will be supplanted by more flexible stress-strain models, which perform even better at fitting the data.
@@ 4.4 Reynolds Stresses
akin to equations (8)-(9) of Torkelsson et al. (2000) but with the additional forcing term. Here, we work with the velocity components $\mathbf{v}=\mathbf{u}-\mathbf{u}_{\rm K}$, relative to the underlying Keplerian flow. The horizontal averaging extracts the coherent sloshing motion, which we will denote $\mathbf{v}_{\rm w}\equiv\langle\rho\mathbf{v}\rangle_{xy}/\bar{\rho}=\{\mathbf{v}\}_{xy}$, where the density average $\bar{\rho}=\langle\rho\rangle_{xy}$. Note that we can identify the in-plane, averaged components with the sloshing amplitudes $A$ and $B$ as defined in equations (6)-(7). Therefore, we decompose the non-Keplerian velocity into a sloshing part and a turbulent residual $\mathbf{u}^{\prime}$ such that $\mathbf{v}=\mathbf{v}_{\rm w}+\mathbf{u}^{\prime}$, where $\{\mathbf{u}^{\prime}\}_{xy}=0$. Propagating this splitting through equations (25)-(26), we find
@@
Figure 9: The rate of strain and stress relationship for our fiducial hydrodynamic run. The upper panels show the phase-folded, vertical profiles for: (a) $v_{x,{\rm w}}$, (b) $v_{y,{\rm w}}$, (c) $\mathcal{R}_{xz}$, and (d) $\mathcal{R}_{yz}$. In panel (e), we extract the best-fitting, sloshing amplitudes $A(t)$ (black) and $B(t)$ (red) across the orbital phase. In panel (f), we similarly extract the best-fitting stress amplitudes, $R_{xz}(t)$ (black-dashed) and $R_{yz}(t)$ (red-dashed).
In Section 4.3, we examined the simplest viscous closure for setting the equilibrium sloshing behavior. Now we are able to directly interrogate the relationship between these averaged flows and the hydrodynamic stress to educate an improved closure model for the turbulence. In Fig. 9, we compare the sloshing and stress response. In the upper panels (a) and (b), we plot $v_{x,{\rm w}}$ and $v_{y,{\rm w}}$ as a function of $z$ over an orbital forcing period, where we have phase-folded in the profiles from all snapshots during the interval $40<t/T<60$. Similarly, we perform this procedure with the Reynolds stresses, $\mathcal{R}_{xy}$ and $\mathcal{R}_{yz}$, which are shown in panels (c) and (d), respectively. The sinusoidal profile for the bulk sloshing motion recapitulates the findings seen in panels (b) and (c) of Fig. 7. Meanwhile, we see that the Reynolds stresses develop an obvious cosine structure, with extrema collocated with the maximally shearing layers. Indeed, the vertical derivative of this Reynolds stress ensures that the overall spatial dependence cancels in equations (27)-(28). At each of the orbital phases, we can fit a sine function to the sloshing motion to extract $A(t)$ and $B(t)$, akin to panel (d) in Fig. 7, but now for our phase-folded data. These are shown in panel (e) of Fig. 9. Similarly, we fit the Reynolds stresses according to $\mathcal{R}_{ij}=R_{ij}(t)\cos(k_{\rm b}z)$, which are shown in panel (f). Fourier analyzing these curves, we extract the associated amplitudes and phases of the dominant $\Omega_{0}$ frequency component. In keeping with our previous results, $|A|=0.185$ and $\arg(A)=-0.76$ whilst $|B|=0.054$ and $\arg(B)=2.49$. [^5] By linearity, the Fourier series for a signal over many cycles is equal to the phase-folded, average signal over a single cycle. Meanwhile, we find that $|R_{xz}|=0.0178$ and $\arg(R_{xz})=2.48$, whilst $|R_{yz}|=0.0177$ and $\arg(R_{yz})=-1.26$. For the previously assumed, purely viscous model, one typically equates $\mathcal{R}_{ij}=\bar{\rho}\nu_{ij}\partial v_{i,{\rm w}}/\partial z$. Inserting the ansatz for the stress-slosh spatial structure yields
@@
This model suggests that the Reynolds stresses should be $\pi$ out of phase with the sloshing flows. In fact, we see that $\arg(R_{xz})-\arg(A)=1.03\pi$ and $\arg(R_{yz})-\arg(B)=-1.19\pi$. Although these are relatively consistent with the viscous picture, there are clear discrepancies. This reemphasizes the results of Section 4.3, wherein the best viscous fit for the equilibrium sloshing could readily capture the sloshing amplitudes, but not their phase relationship. Therefore, we are encouraged to relax the purely viscous prescription and instead appeal to a more flexible closure model for the turbulence. Indeed, by assuming that $\nu$ is real, we are implicitly stating that the stress responds instantaneously to the rate of strain. More realistically, the turbulent transport depends on the history of the shearing motion – in other words, there is some relaxation or delay timescale between the shear and the stress response. Such time delays are also reminiscent of local simulations of the MRI, wherein small lags are found between energy injection rates, via the boundary stresses, and the thermalization rate, as channel flows take some finite time to subsequently decay via parasitic instabilities (see Sano & Inutsuka, 2001; Gardiner & Stone, 2005). This suggests the use of a viscoelastic model for which the stress and strain rate can have an arbitrary phase relationship (e.g. Ogilvie & Proctor, 2003; Ogilvie & Lesur, 2012). We therefore promote $\nu$ to being a complex quantity. Solving equation (30) for the viscoelastic components gives $\nu_{xz}=0.015+0.001i=0.015e^{0.09i}$ and $\nu_{yz}=0.043-0.030i=0.053e^{-0.60i}$. With this complex viscosity in hand, we can also generalize the predictions of the steady-state, laminar sloshing given by equations (15)-(16). Inserting our new results for $\nu_{xz}$ and $\nu_{yz}$ we find that $|A|_{\rm vis}=0.186$ and $|B|_{\rm vis}=0.054$ with a phase difference of $\arg(B_{\rm vis})-\arg(A_{\rm vis})=1.03\pi$, almost in perfect agreement with the values extracted from the simulation. This should be expected given the added flexibility of the viscoelastic model, where each coefficient $\nu_{iz}$ provides two degrees of freedom, which can then fit two amplitudes and two phases.
@@ 5 MHD results
Having investigated the steady sloshing response for the hydrodynamic case, we now wish to extend our setup to include magnetic fields. This will facilitate a complicated interplay between the MRI and PI, which jointly determines the final sloshing state equilibrium. In Section 5.1, we will characterize the base MRI state before the addition of sloshing flows. Then, in Section 5.2, we will examine how the growth of the PI proceeds in the presence of MRI turbulence. Next, in Section 5.3, we will investigate the resulting equilibrium sloshing state and quantify the vertical magnetic and hydrodynamic stresses at play. Finally, in Section 5.4, we will interrogate the dependence of our results on the forcing amplitude, comparing both the hydrodynamic and magnetized regimes.
@@ 5.1 Characterizing the Base MRI
where $B_{0}=\sqrt{2p_{0}/\beta_{\rm m}^{2}}$ and the plasma-beta parameter is chosen to be $\beta_{\rm m}=400$. [^6] Future experimentation should investigate the effect of different MRI field geometries. Furthermore, since it is known that the saturation of the MRI depends on the viscous and ohmic diffusivity values, we explicitly set these to be $\nu=3.2\times 10^{-4}$ and $\eta=8.0\times 10^{-5}$, giving a magnetic Prandtl number $\textrm{Pd}\equiv\nu/\eta=4$, in keeping with the fiducial setup examined by Fromang et al. (2007). As per our hydrodynamic setup, we perform these simulations on a grid with $(L_{x},L_{y},L_{z})=(9.6672,\pi,1)$ and $(N_{x},N_{y},N_{z})=(1232,200,128)$. We seed the instability with white-noise, relative pressure perturbations of amplitude $0.01$, and then run the simulation for $100T$ – tracking the initial growth, nonlinear saturation, and final quasi-steady state for the MRI.
Figure 10: Characterization of our zero-net-flux, base MRI state. Panel (a): plots $\log_{10}(\beta_{\rm m})$ through the meridional slice $y=0$, showing field amplification and fine turbulent structure. Panel (b): growth and saturation of the horizontal Maxwell stresses $\alpha_{\rm M}$ (red), Reynolds stresses $\alpha_{\rm R}$ (blue) and net stress (black), all normalized by the box-averaged pressure. Panel (c): the density-weighted probability density function ($f_{\rm PDF}$) for the velocity fluctuations $u_{x}^{\prime}$ (black), $u_{y}^{\prime}$ (red) and $u_{z}^{\prime}$ (blue), averaged over simulation snapshots between $40<t/T<100$ in intervals of 10.
Representative summary diagnostics, characterizing the base MRI state, are shown in Fig. 10. In panel (a), we plot $\log_{10}(\beta_{\rm m})$ through the $y=0$ slice at $t=100T$, once the MRI has saturated. Here, we clearly observe the presence of turbulent structure over a range of scales. Furthermore, amplification of the magnetic field leads to eddies in which $\beta_{\rm m}$ is of order unity. The strength of the MRI is typically quantified in terms of the nondimensionalized, horizontal accretion stresses it induces. These consist of the hydrodynamic Reynolds stresses
@@
Here, we have normalized the stresses against the box-averaged pressure $\langle\rho c^{2}\rangle=\rho_{0}c^{2}=p_{0}=1$ and also the shear coefficient $q$, so that $\alpha$ can be identified with the Shakura-Sunyaev viscosity parameter acting on the horizontal shear flow. These average stresses are plotted as a function of time in panel (b), where $\alpha_{\rm R}$ (red line) and $\alpha_{\rm M}$ (blue line) are combined into the net stress (black line). Here we see a steep initial rise phase as the MRI grows over tens of dynamical orbits. As the instability saturates, the stresses turn over before leveling off beyond $t=40T$. The stresses are averaged over the gray shaded interval between $40<t/T<100$, yielding typical values for $\alpha_{\rm M}=0.0069$ and $\alpha_{\rm R}=0.0014$. As per usual, the Maxwell stresses dominate over the hydrodynamic effects. In the final panel (c), we plot the density-weighted, normalized histogram ($f_{\rm PDF}$) for the velocity fluctuations, averaged over snapshots at $t/T\in[40,100]$ in intervals of 10. Here we see that the horizontal fluctuations tend to have a wider variance, with characteristic standard deviations $\sigma_{x}=0.085$ and $\sigma_{y}=0.079$, versus $\sigma_{z}=0.056$ for the vertical velocity. For a strong sloshing signal to visibly emerge in the forced simulations, it should manifest on velocities larger than these. However, even small sloshing amplitudes, disguised by the turbulence, can be extracted in an average sense.